OpenStudy (anonymous):

what is the slope of 9x-4xlny=3 at (1/3,1)??

OpenStudy (slaaibak):

use implicit differentiation. 9 - 4lny - 4x/y (dy/dx) = 0 (9 - 4ln y) / (4x/y) = dy/dx Now plug in (1/3, 1)

OpenStudy (anonymous):

thanks man